Determine the value of a so that the average rate of change of the function h(x)= x^2+3x+2 on the interval -3<=x<=a is -1.
Answer by htmentor(1343)

Average rate of change = (h(x2) - h(x1))/(x2 - x1), where the interval is [x1,x2]
-1 = a^2 + 3a + 2 - ((-3)^2 - 3*3 + 2)/(a - -3)
-1 = (a^2 + 3a)/(a + 3) = a(a + 3)/(a + 3) = a
Ans: a = -1